# Prospect of GRB-Neutrino Detection with Enhanced Neutrino Detectors

### v1

@@ Abstract
Gamma-ray bursts (GRBs) have long been proposed as a potential source of high-energy neutrinos. Although no confirmed association between GRBs and neutrinos has been established, meaningful constraints have been placed on GRB prompt emission models. The non-detection of neutrinos, reported by the IceCube collaboration, from both single and stacked GRB events suggests that the radiation zone is likely located at a considerable distance from the central engine, where the photon number density is relatively low. Here, we estimate future GRB-neutrino detection probabilities with more sensitive detectors than IceCube and explore the constraints on models if GRB neutrinos remain undetected despite improved sensitivity. Our findings reveal that if the effective area of a future neutrino detector can be enhanced by a factor of 10 compared to IceCube IC86-II, there is a high likelihood of detecting neutrinos from a GRB 221009A-like event, even in the context of the ICMART model, which exhibits the lowest efficiency in neutrino production. With such an advanced detector (enhanced by a factor of 10) and $5$ to $10$ years of data accumulation, neutrinos from stacked GRBs should be identifiable, or several popular models for GRB prompt emission (e.g., the dissipative photosphere model and internal shock model) could be effectively ruled out.
Gamma-ray bursts (629)
@@ 1 Introduction
In the radiation zone of gamma-ray bursts (GRBs), a large number of gamma-ray photons are generated, either through thermal or non-thermal processes. Simultaneously, protons within the relativistic GRB jet are expected to be accelerated and acquire a relativistic random motion. Consequently, interactions between protons and gamma-ray photons, known as $p\gamma$ interactions, are unavoidable, leading to the production of high-energy neutrinos (e.g. Zhang, 2018). This, as well as the interactions between baryons, make GRB events a compelling candidate as a potential source of high-energy neutrinos (Waxman & Bahcall, 1997; Dermer & Atoyan, 2003; Razzaque et al., 2003; Guetta et al., 2004; Murase et al., 2006a; Murase & Nagataki, 2006; Hümmer et al., 2012; Murase et al., 2013; Bustamante et al., 2015; Tamborra & Ando, 2015; Biehl, D. et al., 2018; Pitik et al., 2021; Ai & Gao, 2023; Rudolph et al., 2023).
@@
However, even for GRB 221009A, the brightest of all time burst, no associated neutrinos were detected by IceCube (Aiello et al., 2024) . The IceCube collaboration reported an upper limit on the neutrino flux associated with this event (IceCube Collaboration, 2022), leading to constraints on GRB models (Murase et al., 2022; Ai & Gao, 2023; Guarini et al., 2023; Rudolph et al., 2023; Veres et al., 2024). It is likely that the radiation radius of GRB 221009A is relatively far from the central engine, which is consistent with the ICMART model, featuring a jet dominated by magnetic fields (Yang et al., 2023). The internal shock model remains viable if the jet has a relatively high bulk Lorentz factor (Murase et al., 2022; Ai & Gao, 2023), whereas the dissipative photosphere model is disfavored (Ai & Gao, 2023).
@@
Considering the difficulty of detecting neutrinos from a single source, stacked neutrino detection across multiple sources has become a more popular approach. As early as 2011, Abbasi et al. (2011) searched for neutrino emission within a one-day time window before and after GRBs from 2008 to 2010. Later, in 2015, Aartsen et al. (2015) analyzed the prompt emission phases of 506 GRBs from 2008 to 2012, placing constraints on the parameter space of the fireball model and estimating that GRBs could account for about 1% of the total astrophysical neutrino flux. Then, in 2017, Aartsen et al. (2017b) extended the data range by three years, including 1,172 GRBs from 2008 to 2015. In 2022, Abbasi et al. (2022) further extended the search window to 14 days before or after GRBs to account for the contribution of afterglows to neutrino production. Also in 2022, Lucarelli, Francesco et al. (2023) used ten years of IceCube data to search for neutrinos during both the GRB prompt and afterglow phases. To date, this stacked detection approach has not yet succeeded in detecting GRB neutrinos.
@@
Several near-future neutrino detectors have been proposed or are already under construction, such as IceCube Gen 2 (Aartsen et al., 2021), Baikal-GVD in Lake Baikal (Avrorin et al., 2011), KM3NeT located in the Mediterranean Sea (Adrián-Martínez et al., 2016) and those in the Pacific Ocean, like P-One (Agostini et al., 2020) and TRIDENT (Ye et al., 2024). At this stage, it is of great interest to assess the prospects for detecting GRB neutrinos as detector sensitivity improves in the future. What is the detection probability for individual bright sources similar to GRB 221009A? After 5 to 10 years of cumulative observations, what is the likelihood of detecting GRB neutrinos through stacked neutrino detection methods? If GRB neutrino detection remains elusive, how constraining would this be for existing models? This work will delve into these questions. To ensure the generality of our research conclusions, we do not focus on the sensitivity of any specific proposed detector but instead calculate by increasing the effective area of IceCube IC86-II by a certain factor.
@@ 2.1 GRB prompt emission
The gamma-ray photons produced during the prompt emission of GRBs provide the source for $p\gamma$ interactions. The photon number density, as well as that of other particles, depend on the radius of the emission region, which increase as getting closer to the central engine. The typical emission radius varies with different GRB models, thus makes the predicted neutrino flux model-dependent. In this work, we include three popular GRB models:
  • Dissipative photosphere model: the gamma-ray photons are generated and trapped in the opaque jet until it reaches the photosphere. The radius of the Thomson scattering photosphere can be estiamted as $R_{\text{ph}}\simeq 3.7\times 10^{11}{\rm cm}~{}L_{w,52}\Gamma^{-3}_{2.5}$ (Meszaros et al., 2001) where $L_{w}=L_{\rm GRB}/\epsilon_{e}$ and the convention $Q_{x}=Q/{10^{x}}$ is used hereafter, where Q can be any quantity. Shocks, neutron-proton collisions and magnetic reconnection are supposed to happen below the photosphere, so that protons are also expected to be accelerated (Rees & Mészáros, 2005; Pe’Er et al., 2007; Giannios, 2008; Zhang & Yan, 2010; Rudolph et al., 2023).
  • Internal shock (IS) model: the collision of different layers of a GRB jet would excite internal shocks, which can accelerate particles and emit gamma-ray photons. This could occur beyond the photosphere of gamma-rays. The typical radius can be estimated as $R_{\rm IS}=2\Gamma^{2}c\delta t_{\rm min}/\left(1+z\right)$ (Rees & Meszaros, 1994; Daigne & Mochkovitch, 1998; Zhang, 2018), where $\delta t_{\rm min}$ stands for the minimum variation timescale for the GRB light curve. Using $\delta t_{\rm min}\sim 0.1$ s and $\Gamma$ in order of $10-100$, the photon emission and proton acceleration occur at a typical internal shock radius as about $10^{12}-10^{13}~{}{\rm cm}$.
  • ICMART model: when the GRB jet is Poynting flux dominated, the global magnetic field may remain un-dissipated beyond $R_{\rm IS}$ and $R_{\rm ph}$. Internal collisions help to destroy the ordered magnetic fields, and a strong run-away magnetic dissipation process occurs at a large radius $R_{\rm ICMART}\sim\Gamma^{2}c\delta t_{\text{slow}}\sim 10^{15}\ \text{cm}$ (Bing & Huirong, 2011), where $\delta t_{\text{slow}}\sim 1$ s is the slow variability component in the GRB light curves (Gao et al., 2012). Particles, either directly accelerated in the reconnection zones or stochastically accelerated within the turbulent regions, emit synchrotron radiation photons, which power the observed prompt emission of GRBs (Zhang & Yan, 2010; Zhang & Zhang, 2014; Shao & Gao, 2022).
@@
When high-energy protons interact with photons of proper energy, they would be in $\Delta$-resonance and produce $\Delta^{+}$. Then the $\Delta^{+}$ decays into mesons, which further decay into leptons and neutrinos (Zhang, 2018; Kimura, 2022). The process can be described as
@@
Besides $\Delta$-resonance, direct pion production and multiple-pion production channels could produce $\pi^{+}$, whose cross sectionreaction is only a factor of a few smaller, thus the contributions from them cannot be ignored. When direct pion production and multiple-pion production channels are considered, the portions of producing $\pi^{+}$and $\pi^{0}$ become 1/2 and 1/2, respectively, (Zhang & Kumar, 2013; Zhang, 2018).
The relationship between the luminosity of neutrino emission and the total energy of the protons can be expressed as
[…]
where $n_{\nu}=\frac{dN_{\nu}}{dE_{\nu}}$ is the neutrino number spectrum. $\epsilon_{\nu,\text{min}}$ and $\epsilon_{\nu,\text{max}}$ refer to the minimum and maximum energies of neutrinos, which correspond to the maximum energy ($E_{p,\text{min}}$) and minimum energy ($E_{p,\text{max}}$) of the accelerated protons, respectively, with the relation $\epsilon_{\nu}\sim 0.05E_{p}$. This is because the $\pi^{+}$ meson carries approximately $1/5$ of the proton energy, and each lepton shares $1/4$ of the $\pi^{+}$ energy. The normalization coefficient is derived as $A=f_{p\gamma}\times\frac{1}{2}\times\frac{1}{4}\times f_{\rm cooling}$, where $f_{p\gamma}$ is the pion production efficiency. The ×$\times\frac{1}{2}$ is because there is a probability of $1/2$ to generate $\pi^{+}$. The ×$\times\frac{1}{4}$ is because the energy of a neutrino is about $1/4$ of $\pi^{+}$ (Waxman & Bahcall, 1997; Zhang & Kumar, 2013). $f_{\rm cooling}$ represents the fraction of intermediate products, such as $\pi^{+}$ and $\mu^{+}$, that have cooled before neutrinos are produced (Zhang & Kumar, 2013; Kimura, 2022).
$E_{\rm p,max}$ could be estimated by equaling the dynamical timescale $t^{\prime}\approx R/\Gamma c$ with the acceletating timescale of protons $t^{\prime}_{\rm acc}\approx E^{\prime}_{p}/\left(eB^{\prime}c\right)$, where $\Gamma$ is bulk motion Lorentz factor of the GRB jet. The“$\prime$” denotes physical quantities defined in the jet’s comoving frame, and hereafter. Hence, we have
$E_{p,\text{max}}\leq 7.59\times 10^{11}\text{ GeV}\left(\frac{\epsilon_{B}}{\epsilon_{e}}\right)^{1/2}L_{\text{GRB},52}^{1/2}\Gamma_{2.5}^{-1}$
$E_{p,\text{min}}>\Gamma m_{p}c^{2}=2.56\times 10^{2}\Gamma_{2.5}\text{ GeV}.$ (4)
The proton spectrum can be expressed as a power law that ${dN_{p}}/{dE_{p}}\propto E_{p}^{-s}$ with s $=2$ adopted (Kimura, 2022). Specifically, it is written as
$\frac{dN_{p}}{dE_{p}}$ $=\frac{\left(\epsilon_{p}/\epsilon_{e}\right)E_{\text{GRB}}E_{p}^{-2}}{\ln\left(E_{p,\text{max}}/E_{p,\text{min}}\right)},$ (5)
where $E_{\text{GRB}}$ is the isotropic energy of the GRB. We assume that a fraction $\epsilon_{e}$ of the dissipated energy is transferred to electrons and fully radiated as gamma rays, a fraction $\epsilon_{p}$ is transferred to protons, and a fraction $\epsilon_{B}$ goes into the random magnetic field.
[…]
where $\gamma_{p}$ is the random-motion Lorentz factor of proton in the comving frame and $\bar{\epsilon}_{\gamma}$ is the energy of photo in the proton’s rest frame. $\kappa_{p\gamma}\approx 0.2$ is the inelasticity coefficient of the photon in the proton’s rest frame. $\sigma_{p\gamma}(\bar{\epsilon}_{\gamma})$ is the cross-section of the $p\gamma$ interaction in the proton’s rest frame. Here, we do not treat it as a constant but rather as a function of the photon energy $\epsilon_{\gamma}$, which is shown in appendix A. $n_{\gamma}$ is the spectrum of gamma-ray bursts in the jet’s comving frame, which is assumed to be in the form of a band spectrum (Poolakkil et al., 2021). Then, we can calculate the generation efficiency of pions as (Waxman & Bahcall, 1997; Kimura, 2022)
$f_{p\gamma}\approx 1-\exp(-{t_{p\gamma}^{-1}}/{t_{\rm dyn}^{-1}})\approx{\rm min}({t_{p\gamma}^{-1}}/{t_{\rm dyn}^{-1}},1).$ (7)
The synchrotron cooling of $\pi^{+}$($\mu^{+}$) would have a suppressive effect on the production of neutrinos. The decay timescale of $\pi^{+}$ is $t^{\prime}_{\pi^{+},\rm dec}=2.6\times 10^{-8}\rm{s}~{}\gamma_{\pi^{+}}$ , where $\gamma_{\pi^{+}}$ is the Lorentz factor for $\pi^{+}$ in the jet’s comoving frame. For the relativistic $\pi^{+}$, the synchrotron cooling timescale can be calculated as $t^{\prime}_{\pi,\text{syn}}={6\pi m_{\pi^{+}}c}/{\gamma_{\pi}\sigma_{T,\pi^{+}}B^{\prime 2}}$ where $m_{\pi^{+}}=0.15m_{p}$ is the rest mass of $\pi^{+}$. The Thomson scattering cross section of $\pi^{+}$ can be estimated from that of the electrons as $\sigma_{T,\pi^{+}}=\left(m_{e}/m_{\pi^{+}}\right)^{2}\sigma_{T,e}$. $\gamma_{\pi^{+}}$ corresponds to the energy of neutrino to be generated. Since the energy of $\pi^{+}$ would be shared nearly equally by four leptons, we obtain $E_{\nu}=\frac{1}{4}D\gamma_{\pi}^{+}m_{\pi^{+}}c^{2}$, where $D\approx\Gamma$ is the Doppler factor  [^1] We consider the case that the observational angle is $1/\Gamma$, where the emissivity of GRBs reaches the maximum.. The cooling factor is then written as
$f_{\rm cooling}\approx 1-\exp(-({t_{\rm syn}^{-1}+t_{\rm dyn}^{-1}})/{t_{\rm dec}^{-1}}).$ (8)
Finally, the predicted neutrino fluence can be derived as
@@
where $S_{\gamma}$ is the gamma ray fluence which we observe.
@@ 3 The detection prospects of GRB 221009A-like events
The gamma-ray burst GRB 221009A is often referred to as the “brightest of all time”. This event was first triggered by the Fermi/Gamma-ray Burst Monitor at 3:16:59 UT on 2022 October 9, with a time-integrated energy flux within the 10–1000 keV range reported as $\left(2.912\pm 0.001\right)\times 10^{-2}\ \text{erg}\ \text{cm}^{-2}$. The peak photon number flux reached $\left(2385\pm 3\right)\ \text{cm}^{-2}\ \text{s}^{-1}$, sustaining this level for $1.024$ s. Using measurements from the Swift observatory, this event was localized at right ascension $\alpha=288.2645^{\circ}$ and declination $\delta=+19.7735^{\circ}$ (Dichiara et al., 2022), with a host galaxy identified at a redshift of $z=0.151$. Consequently, its isotropic energy reaches $\sim 10^{55}\ \text{erg}$ (An et al., 2023; Yang et al., 2023; Lan et al., 2023), making it a highly promising candidate to exhibit a neutrino counterpart, although it was not detected (Abbasi et al., 2023).
@@
where $F_{\nu}=\phi_{\nu}/(E_{\nu}^{2}T)$ is the specific number flux of neutrinos, with $T$ as the duration of observation length. $A_{\rm eff}$ is the effective areas of the neutrino detector. Here we use IceCube IC86-II Effective Area (IceCube Collaboration, 2021), which depends on the neutrino energy and the declination of the source in the sky. The effective areas corresponding to several typical declinations as a function of neutrino energy are shown in Figure 1.
@@
The predicted neutrino fluence associated with GRB 221009A from the dissipative photosphere, internal shock, and ICMART models, along with the corresponding $90\%$ upper limit under the non-detection condition with IceCube, is shown in Figure 2. Here, we adopt $\epsilon_{p}/\epsilon_{e}=3$ and $\epsilon_{B}/\epsilon_{e}=1$ for all models. For the internal shock model, $\delta t_{\text{min}}=0.01\,\text{s}$, and for the ICMART model, $R_{\rm ICMART}=10^{15}$ cm. With the predicted fluence, we can get the corresponding neutrino number to be detected by IceCube, that is $N_{\rm ph}=13.0132$, $N_{\rm IS}=3.5410$, and $N_{\rm ICMART}=0.2109$. Thus, for each model, the detection probability can be calculated as $P_{\rm ph}=99.99\%$, $P_{\rm IS}=97.10\%$, and $P_{\rm ICMART}=19.02\%$.
Figure 2: The solid lines represent the predicted neutrino spectrum for GRB 221009A based on the internal shock, dissipative photosphere, and ICMART models, respectively. The indices for the Band function are fitted as $\alpha=0.97$ and $\beta=2.37$. The isotropic energy, $E_{\text{GRB}}=1.15\times 10^{55}\ \text{erg}$, the isotropic luminosity, $L_{\text{GRB}}=1.9\times 10^{52}\ \text{erg}\ \text{s}^{-1}$ and the $\Gamma=300$ are adopted. For all models, $\epsilon_{B}/\epsilon_{e}=1$ and $\epsilon_{p}/\epsilon_{e}=3$ are adopted. For the internal shock model, $\delta t_{\text{min}}=0.01\ \text{s}$ is adopted. For ICMART model, $R_{\rm ICMART}=10^{15}$ cm is adopted. The dashed lines represent the $90\%$ confidence-level upper limit of the fluence under non-detection conditions with IceCube, with effective areas of IceCube IC86-II applied.
@@
Despite GRB 221009A being referred to as a “once-in-a-millennium” GRB, it is not particularly unique. Lan et al. (2023) suggests that it is simply an ordinary nearby GRB with extraordinary observational properties. Therefore, in the future, if a GRB 221009A-like event occurs at different sky positions where the neutrino detector can achieve a larger effective area, neutrinos from such events may have chance to be detected.
@@
Figure 3: The prospect to detect a GRB 221009A-like event. The horizontal axis represents redshift of the event, and the vertical axis represents the magnification factor relative to the effective area of IceCube IC86-II. The black line represents the redshift of the GRB 221009A. The colored solid lines represent the detection scenarios for different models at the current declination of GRB 221009A, while the dashed lines correspond to the scenarios at the declination where the effective area is maximized.

### v2

# Prospect of Gamma-Ray Burst Neutrino Detection with Enhanced Neutrino Detectors

@@ Abstract
Gamma-ray bursts (GRBs) have long been proposed as a potential source of high-energy neutrinos. Although no confirmed association between GRBs and neutrinos has been established, meaningful constraints have been placed on GRB prompt emission models. The nondetection of neutrinos, reported by the IceCube Collaboration, from both single and stacked GRB events suggests that the radiation zone is likely located at a considerable distance from the central engine, where the photon number density is relatively low. Here, we estimate future GRB neutrino detection probabilities using detectors with a higher simulated sensitivity than IceCube and explore the constraints on models if GRB neutrinos remain undetected despite improved sensitivity. Our findings reveal that if the effective area of a future neutrino detector can be enhanced by a factor of 10 compared to IceCube IC86-II, there is a high likelihood of detecting neutrinos from a GRB 221009A-like event, even in the context of the ICMART model, which exhibits the lowest efficiency in neutrino production. With such an advanced detector (enhanced by a factor of 10) and 5–10 yr of data accumulation, neutrinos from stacked GRBs should be identifiable, or several popular models for GRB prompt emission (e.g., the dissipative photosphere model and internal shock model) could be effectively ruled out.
Gamma-ray bursts (629); Cosmological neutrinos (338); Neutrino telescopes (1105)
@@ 1 Introduction
In the radiation zone of gamma-ray bursts (GRBs), a large number of gamma-ray photons are generated, either through thermal or nonthermal processes. Simultaneously, protons within the relativistic GRB jet are expected to be accelerated and acquire a relativistic random motion. Consequently, interactions between protons and gamma-ray photons, known as $p\gamma$ interactions, are unavoidable, leading to the production of high-energy neutrinos (e.g. Zhang, 2018). This, as well as the interactions between baryons, make GRB events a compelling candidate as a potential source of high-energy neutrinos (e.g. Waxman & Bahcall, 1997; Dermer & Atoyan, 2003; Razzaque et al., 2003; Guetta et al., 2004; Murase & Nagataki, 2006; Hümmer et al., 2012; Murase et al., 2013; Bustamante et al., 2015; Tamborra & Ando, 2015; Biehl, D. et al., 2018).
@@
However, even for GRB 221009A, the brightest-of-all-time burst, no associated neutrinos were detected by IceCube (Aiello et al., 2024) . The IceCube Collaboration reported an upper limit on the neutrino flux associated with this event (IceCube Collaboration, 2022), leading to constraints on GRB models (Murase et al., 2022; Ai & Gao, 2023; Rudolph et al., 2023; Veres et al., 2024; Ou et al., 2024). It is likely that the radiation radius of GRB 221009A is relatively far from the central engine, which is consistent with the ICMART model, featuring a jet dominated by magnetic fields (Yang et al., 2023). The internal shock model remains viable if the jet has a relatively high bulk Lorentz factor (Murase et al., 2022; Ai & Gao, 2023), whereas the dissipative photosphere model is disfavored (Ai & Gao, 2023).
@@
Considering the difficulty of detecting neutrinos from a single source, stacked neutrino detection across multiple sources has become a more popular approach. As early as 2011, Abbasi et al. (2011) searched for neutrino emission within a 1 day time window before and after GRBs from 2008 to 2009. Later, in 2015, Aartsen et al. (2015) analyzed the prompt emission phases of 506 GRBs from 2008 to 2012, placing constraints on the parameter space of the three models and estimating that GRBs could account for about 1% of the total astrophysical neutrino flux. Then, in 2017, Aartsen et al. (2017b) extended the data range by 3 yr, including 1172 GRBs from 2008 to 2015. In 2022, Abbasi et al. (2022) further extended the search window to 14 days before or after GRBs to account for the contribution of afterglows to neutrino production. Also in 2022, Lucarelli, Francesco et al. (2023) used 10 yr of IceCube data to search for neutrinos during both the GRB prompt and afterglow phases. To date, this stacked detection approach has not yet succeeded in detecting GRB neutrinos.
@@
Several near-future neutrino detectors have been proposed or are already under construction, such as IceCube Gen2 (Aartsen et al., 2021), Baikal-GVD in Lake Baikal (Avrorin et al., 2011), KM3NeT located in the Mediterranean Sea (Adrián-Martínez et al., 2016) and those in the Pacific Ocean, like P-One (Agostini et al., 2020) and TRIDENT (Ye et al., 2024). At this stage, it is of great interest to assess the prospects for detecting GRB neutrinos as detector sensitivity improves in the future. What is the detection probability for individual bright sources similar to GRB 221009A? After 5-10 yr of cumulative observations, what is the likelihood of detecting GRB neutrinos through stacked neutrino detection methods? If GRB neutrino detection remains elusive, how constraining would this be for existing models? This work will delve into these questions. To ensure the generality of our research conclusions, we do not focus on the sensitivity of any specific proposed detector but instead calculate by increasing the effective area of IceCube IC86-II by a certain factor.
In this work, the CGS unit system is adopted, and the convention $Q_{x}=Q/{10^{x}}$ is used, where $Q$ represents any physical quantity. This convention applies only when $x$ is a number.
@@ 2.1 GRB prompt emission
The gamma-ray photons produced during the prompt emission of GRBs provide the source for $p\gamma$ interactions. The photon number density, as well as that of other particles, depends on the radius of the emission region, which increases as it gets closer to the central engine. The typical emission radius varies with different GRB models, thus makes the predicted neutrino flux model dependent. In this work, we include three popular GRB models:
  • Dissipative photosphere model. The gamma-ray photons are generated and trapped in the opaque jet until it reaches the photosphere. The radius of the Thomson scattering photosphere can be estimated as $R_{\text{ph}}\simeq 3.7\times 10^{11}{\rm cm}~{}L_{w,52}\Gamma^{-3}_{2.5}$ (Meszaros et al., 2001), where $\Gamma$ is the bulk Lorentz factor of the jet and $L_{w}=L_{\rm GRB}/\epsilon_{e}$. Here, we assume that a fraction $\epsilon_{e}$ of the dissipated energy in the jet goes into electrons, which then convert into gamma-ray photons. Shocks, neutron-proton collisions and magnetic reconnection are supposed to happen below the photosphere, so that protons are also expected to be accelerated (Rees & Mészáros, 2005; Pe’Er et al., 2007; Giannios, 2008; Zhang & Yan, 2010; Rudolph et al., 2023).
  • Internal shock (IS) model. The collision of different layers of a GRB jet would excite internal shocks, which can accelerate particles and emit gamma-ray photons. This could occur beyond the photosphere of gamma rays. The typical radius can be estimated as $R_{\rm IS}=2\Gamma^{2}c\delta t_{\rm min}/\left(1+z\right)$ (Rees & Meszaros, 1994; Daigne & Mochkovitch, 1998; Zhang, 2018), where $\delta t_{\rm min}$ stands for the minimum variation timescale for the GRB light curve. Using $\delta t_{\rm min}\sim 0.1$ s and $\Gamma$ in order of 10-100, the photon emission and proton acceleration occur at a typical internal shock radius as about $10^{12}-10^{13}~{}{\rm cm}$.
  • ICMART model. When the GRB jet is Poynting-flux dominated, the global magnetic field may remain undissipated beyond $R_{\rm IS}$ and $R_{\rm ph}$. Internal collisions help to destroy the ordered magnetic fields, and a strong runaway magnetic dissipation process occurs at a large radius $R_{\rm ICMART}\sim\Gamma^{2}c\delta t_{\text{slow}}\sim 10^{15}\ \text{cm}$ (Zhang & Yan, 2010), where $\delta t_{\text{slow}}\sim 1$ s is the slow variability component in the GRB light curves (Gao et al., 2012). Particles, either directly accelerated in the reconnection zones or stochastically accelerated within the turbulent regions, emit synchrotron radiation photons, which power the observed prompt emission of GRBs (Zhang & Yan, 2010; Zhang & Zhang, 2014; Shao & Gao, 2022).
@@
When high-energy protons interact with photons of proper energy, they would be in $\Delta$ resonance and produce $\Delta^{+}$. Then, the $\Delta^{+}$ decays into mesons, which further decay into leptons and neutrinos (e.g. Mücke et al., 1999). The process can be described as
@@
Besides $\Delta$-resonance, direct pion production and multiple-pion production channels can produce $\pi^{+}$, whose cross-section reaction is only a factor of a few smaller; thus, the contributions from them cannot be ignored. When direct pion production and multiple-pion production channels are considered, the portions of producing $\pi^{+}$and $\pi^{0}$ become 1/2 and 1/2, respectively (Zhang & Kumar, 2013; Zhang, 2018).
We follow Waxman & Bahcall (1997) and Kimura (2022) to calculate the predicted neutrino fluence, which can be written as
@@
where $S_{\gamma}$ is the gamma ray fluence that we observe. $E_{p,\text{max}}$ and $E_{p,\text{min}}$ are the maximum energy and minimum energy of the accelerated protons. Here, we assume that a fraction $\epsilon_{p}$ of the dissipated energy in the jet goes into protons, while another fraction $\epsilon_{B}$ goes into the magnetic field. Together with the fraction previously mentioned that goes into electrons, the dissipated energy is thus divided into three components in total. $f_{p\gamma}$ is the pion production efficiency. $f_{\rm cooling}$ represents the fraction of intermediate products, such as $\pi^{+}$ and $\mu^{+}$, that have cooled before neutrinos are produced. The detailed derivations of the above can be found in the Appendix A.
@@ 3 The detection prospects of GRB 221009A-like events
The gamma-ray burst GRB 221009A is often referred to as the “brightest of all time” (Burns et al., 2023; Lesage et al., 2023; Williams et al., 2023; An et al., 2023; Frederiks et al., 2023). This event was first triggered by the Fermi/Gamma-ray Burst Monitor at 3:16:59 UT on 2022 October 9, with a time-integrated energy flux within the 10–1000 keV range reported as $\left(2.912\pm 0.001\right)\times 10^{-2}\ \text{erg}\ \text{cm}^{-2}$. The peak photon number flux reached $\left(2385\pm 3\right)\ \text{cm}^{-2}\ \text{s}^{-1}$, sustaining this level for $1.024$ s (Lesage et al., 2022). Using measurements from the Swift observatory, this event was localized at right ascension $\alpha=288.2645^{\circ}$ and declination $\delta=+19.7735^{\circ}$ (Dichiara et al., 2022), with a host galaxy identified at a redshift of $z=0.151$. Consequently, its isotropic energy reaches $\sim 10^{55}\ \text{erg}$ (An et al., 2023; Yang et al., 2023; Lan et al., 2023), making it a highly promising candidate to exhibit a neutrino counterpart that can be seen by IceCube, although it was not detected (Abbasi et al., 2023).
@@
where $F_{\nu}=\phi_{\nu}/(E_{\nu}^{2}T)$ is the specific number flux of neutrinos, with $T$ as the duration of observation length. $A_{\rm eff}$ is the effective areas of the neutrino detector. Here, we use IceCube IC86-II effective area (IceCube Collaboration, 2021), which depends on the neutrino energy and the decl of the source in the sky. The effective areas corresponding to several typical declinations as a function of neutrino energy are shown in Figure 1.
@@
The predicted neutrino fluence associated with GRB 221009A from the dissipative photosphere, internal shock, and ICMART models, along with the corresponding $90\%$ upper limit under the nondetection condition with IceCube, is shown in Figure 2. Here, we adopt $\epsilon_{p}/\epsilon_{e}=3$ and $\epsilon_{B}/\epsilon_{e}=1$ for all models. For the internal shock model, $\delta t_{\text{min}}=0.01\,\text{s}$, and for the ICMART model, $R_{\rm ICMART}=10^{15}$ cm. With the predicted fluence, we can get the corresponding neutrino number to be detected by IceCube, that is, $N_{\rm ph}=13.0132$, $N_{\rm IS}=3.5410$, and $N_{\rm ICMART}=0.2109$. Thus, for each model, the detection probability can be calculated as $P_{\rm ph}=99.99\%$, $P_{\rm IS}=97.10\%$, and $P_{\rm ICMART}=19.02\%$.
Figure 2: The solid lines represent the predicted neutrino spectrum for GRB 221009A based on the internal shock, dissipative photosphere, and ICMART models, respectively. The indices for the Band function are fitted as $\alpha=0.97$ and $\beta=2.37$. The isotropic energy, $E_{\text{GRB}}=1.15\times 10^{55}\ \text{erg}$, the isotropic luminosity, $L_{\text{GRB}}=1.9\times 10^{52}\ \text{erg}\ \text{s}^{-1}$ and the $\Gamma=300$ are adopted. For all models, $\epsilon_{B}/\epsilon_{e}=1$ and $\epsilon_{p}/\epsilon_{e}=3$ are adopted. For the internal shock model, $\delta t_{\text{min}}=0.01\ \text{s}$ is adopted. For ICMART model, $R_{\rm ICMART}=10^{15}$ cm is adopted. The dashed lines represent the $90\%$ confidence-level upper limit of the fluence under nondetection conditions with IceCube, with effective areas of IceCube IC86-II applied.
@@
Despite the unique characteristics of GRB 221009A, primarily its exceptionally bright luminosity, studies like Lan et al. (2023) suggest that it is a nearby bright GRB with properties similar to those of energetic GRBs. Therefore, in the future, if a GRB 221009A-like event occurs at different sky positions where the neutrino detector can achieve a larger effective area, neutrinos from such events may have chance to be detected.
@@
Figure 3: The prospect to detect a GRB 221009A-like event. The horizontal axis represents the redshift of the event, and the vertical axis represents the magnification factor relative to the effective area of IceCube IC86-II. The black line represents the redshift of the GRB 221009A. The colored solid lines represent the detection limits for different models at the current decl of GRB 221009A, while the dashed lines correspond to scenarios in which GRB 221009A is located near the equator, where the effective area is maximized. The effective area is calculated by integrating over the neutrino energy range from $10^{2}$eV to $10^{9}$ GeV. The arrows indicate the parameter spaces in which the source is detectable.